You just won the $20,000,000 lottery price. Either you can take the annual payments of $1,000,000 per year for 20 years or you can elect to receive a lump sum value of the annual payments using a discount rate of 6%. You have elected to receive the lump sum amount. What is the net lump sum amount you will received after applying Federal and State Income Taxes; 25% and 12% respectively. What is the Net Lump Sum distributed after taxes?

A. $11, 469,921

B. $4,243,871

C. $7,226,050
Answer #1
Winnings $       20,000,000
PV @ 6% 0.94340
Lump Sum Pv @ $       18,867,925
Federal Tax @ 25% $         4,716,981
State Tax @ 12% $         2,264,151
Net Lump Sum $       11,886,792
Winnings 20000000
PV @ 6% =1/(1+6%)^1
Lump Sum Pv @ =C2*C3
Federal Tax @ 25% =C4*25%
State Tax @ 12% =C4*12%
Net Lump Sum =C4-C5-C6